What is the area of A 30 60 90 triangle with a short leg measuring 6 units
Answer by jim_thompson5910(35256)

Short Leg = SL = 6 units

Long leg = LL = 6*sqrt(3) units

Area = (b*h)/2 = (SL*LL)/2 = (6*6*sqrt(3))/2 = 18*sqrt(3) square units

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------